OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f/g (−2).

OpenStudy (anonymous):

\[\large \frac{f}{g}(x)=\frac{f(x)}{g(x)} \]

OpenStudy (anonymous):

f/g(2)=f(2)/g(2) so answer is ((-2)*-3-3)/(2*(-2)+1) = (3)/(-3)=-1